Conseco, Inc., has a debt ratio of 0.43. What are the company's debt-to-equity ratio and equity multiplier?
Answer #1

debt = 0.43

debt + equity =1

equity = 0.57

debttoequity ratio = 0.43/0.57 = 0.75

equity multiplier = 1/equity = 1/0.57 = 1.75